OpenStudy (kirbykirby):

By using both sides of the identity \[\frac{1}{1-x}=\sum_{n=0}^{\infty}x^n\], find the sum of the series \[\sum_{n=0}^{\infty}\frac{4(-1)^n}{2n+1}\]

OpenStudy (kirbykirby):

I get to the point where I have \[\frac{1}{2}(\ln(1+x)-\ln(1-x))+C=\sum_{n=0}^{\infty}\frac{x^{2n+1}}{2n+1}\] which seems close enough to where I need to go? but I don't know what to do next

OpenStudy (zarkon):

look at the arctangent

OpenStudy (zarkon):

\[\tan^{-1}(x)=\sum_{n=0}^{\infty}(-1)^n\frac{x^{2n+1}}{2n+1}\]

OpenStudy (zarkon):

\[\frac{d}{dx}\tan^{-1}(x)=\frac{1}{1+x^2}=\frac{1}{1-(-x^2)}=\cdots\]

OpenStudy (kirbykirby):

oh wait is it if we substitute x=1, we get rid of the annoying 2n+1 power, so arctan(1)= \[\sum_{n=0}^{\infty}\frac{(-1)^n}{2n+1}\] and so to get the 4, we do 4*arctan(1), and arctan(1) = pi/4, so 4*arctan(1) = pi??

OpenStudy (zarkon):

yes

OpenStudy (kirbykirby):

ahh I see :)! I didn't think of using arctan ._. so thanks!

OpenStudy (zarkon):

no problem